# MMDesign: Multi-Modality Transfer Learning for Generative Protein Design

### v1

@@ Abstract
Protein design involves generating protein sequences based on their corresponding protein backbones. While deep generative models show promise for learning protein design directly from data, the lack of publicly available structure-sequence pairings limits their generalization capabilities. Previous efforts of generative protein design have focused on architectural improvements and pseudo-data augmentation to overcome this bottleneck. To further address this challenge, we propose a novel protein design paradigm called MMDesign, which leverages multi-modality transfer learning. To our knowledge, MMDesign is the first framework that combines a pretrained structural module with a pretrained contextual module, using an auto-encoder (AE) based language model to incorporate prior semantic knowledge of protein sequences. We also introduce a cross-layer cross-modal alignment algorithm to enable the structural module to learn long-term temporal information and ensure consistency between structural and contextual modalities. Experimental results, only training with the small CATH dataset, demonstrate that our MMDesign framework consistently outperforms other baselines on various public test sets. To further assess the biological plausibility of the generated protein sequences and data distribution, we present systematic quantitative analysis techniques that provide interpretability and reveal more about the laws of protein design.
@@
Figure 1: (a) The mainstream advanced deep protein design framework, where the GNN-based structure module represents the structural features, followed by an optional contextual module encoding the sequence information, and the decoder responsible for decoding the generated protein sequences. (b) The proposed novel paradigm of the protein design framework consists of two pretrained modules, i.e., the pretrained graph structure module and the pretrained auto-encoder (AE) contextual module. Additional cross-layer cross-modal alignment aims to enhance feature constraints.
1 Introduction
@@
De novo protein design, also known as protein inverse folding, aims to generate protein sequences (i.e., amino acids) based on the corresponding atomic coordinates of the protein backbones. Protein design is crucial for various applications in bioengineering, including drug design and enzyme discovery, but it remains one of the challenging tasks. Recent years have seen significant progress in protein design, thanks to the emergence of deep learning-based methods. In particular, supervised generative models have enabled more efficient and effective prediction of protein sequences[13, 23, 2, 14, 8, 4]. Deep generative models have the potential to learn protein design rules directly from data, making them a promising alternative to traditional approaches such as physics-based energy functions.
@@
The current mainstream and advanced generative protein design frameworks typically follow an encoder-decoder architecture, as illustrated in Figure 1(a). Firstly, a graph is used to represent the structural information in Euclidean space based on the atomic coordinates of the protein backbone. Subsequently, graph neural networks (GNNs) are responsible for encoding graph-based geometric information into higher dimensional representations using a message-passing mechanism. Finally, the decoder module generates the protein sequences through either autoregressive networks based on self-attention[13, 14] or non-autoregressive networks based on MLPs[5, 6, 32]. However, the lack of massively available parallel datasets of protein structure-sequence pairings has become a major bottleneck for deep protein design models, leading to problems such as under-training or over-fitting. Although billions of protein sequences are available, only a few hundred thousand 3D protein structures have been experimentally determined due to the extremely high costs. Hence, as a multi-modality generative task, it is still challenging to train a practical protein design model based on available low-resource data, especially when compared to other established cross-modal tasks such as vision-language tasks.
To address this limitation, various solutions have emerged: 1) The most common method is to focus on protein design architecture, mainly by enhancing the GNN-based structural module’s encoding capability of protein 3D spatial structure (e.g., introducing a robust geometric constraint network for enhanced equivariance and invariance[13, 14]), or extending the structural module with a temporal module (e.g., RNN or self-attention layers) to further learn relationships between nodes [8], as shown in Figure 1(a). Although only a slight performance improvement, configurations such as equivalent GNNs and temporal modules that source this method have become standard in mainstream protein design models. 2) The most straightforward approach is to expand the dataset by generating large-scale pseudo-structures[8] using advanced protein structure prediction models[15, 3, 18]. Despite data augmentation being widely employed, the data construction process is tedious and time-consuming, and pseudo-data might also introduce potential noise. 3) Another newly emerging framework is to introduce a pretrained structural module, which greatly reduces the reliance on supervised data while maintaining satisfactory performance [32]. Although transfer learning is common in fields such as natural language processing and computer vision, it is not widely used for protein design due to the complexity of structure model pretraining and the limited publicly available related works.
Among these solutions, utilizing pretrained structural networks has become a consensus to overcome the low-resource limit. To further enhance the temporal module that complements the structural module with semantic and temporal representation, it is worth considering injecting the prior language knowledge of protein sequences. However, incorporating prior knowledge of both source and target modalities is intractable for supervised tasks, particularly in protein design where there are no reference works available. Fortunately, a new paradigm in the visual-textual domain [33] allows us to introduce prior knowledge of the target modality in an auto-encoder (AE) manner.
@@
Based on the observations above, we propose a novel paradigm for generative protein design, MMDesign, which leverages multi-modality transfer learning through pretrained structural and contextual modules, as illustrated in Figure 1(b). This framework fully exploits pretraining models to overcome the data bottleneck. For the structural module, we incorporate advanced equivariant network-based pretrained structural representation capabilities. For the contextual module, we replace the previous temporal module and decoder module with an AE model, which can integrate comprehensive prior language knowledge and assume the functions of protein semantics encoding and sequence generation.
Furthermore, we incorporate explicit cross-modal alignment constraints to MMDesign, which have been proven effective in multi-modality domains. For example, [20, 7] employ the distillation loss to align the visual and textual modules. This cross-modal alignment can be viewed as a “consistency” metric to measure the distance between two probability distributions [36, 33]. Maintaining consistency across modalities allows the structural module to be constrained by long-range temporal context, as shown in Figure 1(b). Notably, the AE model serves as an implicit cross-modal bridge between the structural and contextual modality, similar to the impact of “consistency”[33].
MMDesign is trained solely on the small dataset CATH [22, 13], yet extensive experiments demonstrate that MMDesign substantially outperforms the state-of-the-art baselines without relying on large-scale supervised data. However, previous studies have focused primarily on average performance, with little attention paid to explaining the generated results and their biological plausibility. We present the first systematic quantitative and comprehensive analysis techniques to assess the plausibility of the generated protein sequences and the distribution of data. The analysis techniques introduced are new for protein design studies, validating the model’s generalization on unseen and unknown datasets, examining the effect of sequence length on the model, assessing the biological plausibility of the generated data distribution and foldability, and performing internal crucial segment discovery and confusion matrix analysis.
The main contributions are as follows:
  • The proposed MMDesign, a novel protein design framework that includes both pretrained structural knowledge and prior contextual semantic knowledge for protein design. This provides a new way of thinking to overcome the bottleneck of other cross-modal protein tasks.
  • The introduction of explicit cross-modal alignment between the structural and contextual modality using a cross-layer consistency constraint and implicit cross-modal alignment using favourable properties of the AE mechanism.
  • Despite only training on a small dataset, MMDesign substantially outperforms other baselines on multiple benchmarks and even surpasses the strong baselines trained on large datasets.
  • More comprehensive interpretable analyses are presented, which provide multidimensional validation and reveal the underlying protein design patterns.
  • The code and trained models will be publicly released.
2 Related Work
@@
The research of protein engineering based on deep learning has attracted more and more attention [12, 11, 10, 19, 9]. Early work on protein design primarily relied on using an energy function that directly models the physical basis of the protein folding state [1]. In recent years, deep learning-based approaches have gained popularity, and generative models that predict the sequences of structures have gradually replaced traditional methods, yielding superior results. For instance, Structured Transformer[13] introduced the formulation of fixed-backbone design as a conditional sequence generation problem. It used invariant features with GNNs and modeled each amino acid as a graph node with edges connecting spatially adjacent amino acids. GVP-GNN [14] further improved GNNs by developing architectures with translation- and rotation-equivariance to enable geometric reasoning, highlighting the powerful structural representation of the geometric vector perceptron (GVP) and GNN module. More recently, ProteinMPNN [4] achieved higher-level sequence recovery on the native backbone by incorporating noise during training and it also improved inference efficiency by performing large-scale side-chain packing calculations. Then PiFold [6] contained a novel residue featurizer and GNN layers to predict protein sequences in a one-shot way. While PiFold improves efficiency and optimizes recovery in single and short chains, it struggles with longer sequences. ESM-IF[8] utilized GVP-based encoder layers to extract geometric features followed by a Transformer. It also employed a pseudo-data augmentation strategy that leveraged the structures of 12 million protein sequences predicted by AlphaFold2[15], achieving an impressive native sequence recovery. However, scaling up the training data by nearly three orders of magnitude significantly increases the training difficulty. Moreover, its better performance does not extend to single and short chains. Overall, the existing works still exhibit limited generalization capabilities without relying on large-scale data. Furthermore, they do not offer a comprehensive biological plausibility analysis of the predicted protein sequences and lack interpretability.
3 Methods
[…]
@@ 3 Methods
The overall training pipeline is divided into two main steps. Step 1 is to initialize the GVPConv-based structural module and pretrain the AE network for the contextual module. The parameters of the structural module are directly derived from the off-the-shelf protein structure model [32], which is trained in a self-supervised manner. The contextual module pertaining can be regarded as a sequence-to-sequence pseudo-translation task. And step 2 is to transfer the pretrained structural module and contextual module into the MMDesign framework. To encourage multi-modality feature interaction, we also introduce cross-layer cross-modal constraints.
3.1 Pretraining Stage
3.1.1 Geometric Structural Module Initialization
@@
The GVPConv structural module is a key component of our MMDesign framework. GVP has been widely used in protein design[14, 8] due to its lightweight and efficient linear transformation, as well as its ability to learn rotation-equivariant transformations of vector features and rotation-invariant transformations of scalar features. In our framework, we use a variant GVP as in ESM-IF [8], which transforms the vector features into local reference frames defined for each amino acid to derive rotation-invariant features.
To ensure that the predicted sequence is independent of the reference frame of the structural coordinates, we enforce invariance under any rotation and translation $T$ of the input coordinates $X$, i.e., $p(Y|X)=p(Y|TX)$, where $Y$ is the predicted sequence. The full graph convolution following GVP utilizes message passing to update the node embeddings in each propagation step, taking into account messages from the neighbor nodes and edges [8, 14].
To obtain a better representation of the protein structures, we initialize the GVPConv parameters as [32], where the GVPConv module is pretrained with guidance from a large-scale protein language model through contrastive alignment constraints.
3.1.2 Contextual Transformer Pretraining
Our AE model is a generic auto-regressive encoder-decoder Transformer architecture[27]. The only modification is the use of learned positional embeddings instead of classical sinusoidal positional embeddings. Shared token embedding and shared positional encoding are used since the encoder and decoder share the same amino acid tokens in the AE model.
@@ 3.1.2 Contextual Transformer Pretraining
To pretrain the AE for the proposed MMDesign framework solely based on the sequence data from the CATH training set (excluding the validation set and test set), we employ a sequence-to-sequence recovery task as a pseudo-translation process in natural language, as shown in Figure 2. This enables the AE to learn contextual semantic knowledge. And cross-entropy loss is used to fit the probability distribution of the generated sequences. Formally, we assume that the input protein sequence of the encoder is denoted as $S_{\text{in}}=\{s_{1},s_{2},\cdots,s_{n}\}$ with $n$ amino acids, and the generated protein sequence of the decoder is denoted as $S_{\text{out}}=\{s^{\prime}_{1},s^{\prime}_{2},\cdots,s^{\prime}_{n}\}$, which corresponds to the probability distribution denoted as $S_{\text{logits}}$. The reference native sequence is $S_{\text{native}}=S_{\text{in}}=\{s_{1},s_{2},\cdots,s_{n}\}$. The sequence cross-entropy loss is as follows:
@@
Since a relatively small training set of only 18k instances is used, the AE network could converge quickly to obtain better initialization parameters.
3.2 Protein Design Framework Training
3.2.1 MMDesign Training Pipeline
In the MMDesign framework, the pretrained structural module and the pretrained contextual module are sequentially concatenated, as illustrated in the lower section of Figure 2. The pretrained contextual module only preserves the encoder and decoder layers of the AE model in step 1 and discards the embedding layers, acting as the repository of prior language knowledge.
[…]
Firstly, the protein backbone $X$ is input to the pretrained GVPConv structural module to obtain the geometric structural features $Z_{\text{struc}}=\{z_{0},z_{1},\cdots,z_{n}\}$ with a dimension $d$ as:
$Z_{\text{struc}}=\text{GVP}(X)\in\mathbb{R}^{n\times d}.$ (2)
Then, the structural features $Z_{\text{struc}}$ are fed to the encoder-decoder layers of the pretrained contextual module sequentially. The features obtained by the encoder are referred to contextual features $Z_{\text{seq}}=\{z^{\prime}_{0},z^{\prime}_{1},\cdots,z^{\prime}_{n}\}$ as:
$Z_{\text{seq}}=Encoder(Z_{\text{struc}})\in\mathbb{R}^{n\times d}.$ (3)
Finally, the decoder decodes the contextual features to generate the output distribution $D_{\text{logits}}=\{o_{0},o_{1},\cdots,o_{n}\}$ of the generated sequence as:
$\begin{split}D_{\text{logits}}&=Decoder(Z_{\text{seq}})\in\mathbb{R}^{n\times M},\\ Y&=\text{argmax}\big{(}\text{Softmax}(D_{\text{logits}})\big{)}\in\mathbb{R}^{n},\end{split}$ (4)
where $M$=20 denotes 20 types of the most common amino acids.
# Models Perplexity Recovery (%)
@@ 3.2.1 MMDesign Training Pipeline
Table 1: Perplexity and recovery scores for MMDesign and other baselines on the CATH, Ts50, and Ts500 datasets. The best results are bolded, and the second best underlined. Group 1 & Group 3 are trained on a small-scale CATH training set only, while Group 2 is trained on a big-scale CATH+AlaphaDB training set. The Short and Single-chain are subsets of CATH test set ALL. *: evaluated on CATH 4.3 test set; Others: evaluated on CATH 4.2 test set. (Due to version differences, several baselines used CATH 4.3. However, both versions are sampled in CATH, so the data distribution and conclusions are similar.)
3.2.2 Generation Objective
The cross entropy loss is commonly used in protein design tasks to measure the discrepancy between the probability distribution of the generated amino acid sequences and the reference native sequences. By minimizing the cross entropy, we can obtain an approximate distribution that better matches the target distribution. In our experiments, we observed that emphasizing the salient features of the output distribution can further improve the performance of our model. To achieve this, we propose an exponential cross entropy (expCE) that accentuates the spiking effect of the output distribution. The expCE is calculated as follows:
@@ 3.2.2 Generation Objective
3.2.3 Cross-modal Alignment
Multi-modality research has shown that explicitly enforcing consistency between different modalities can improve cross-modal frameworks [35, 20, 7, 30, 31, 33]. In this work, we view protein design as a cross-modal task, where the structural features $Z_{\text{struc}}$ obtained from the GVPConv module and the contextual features $Z_{\text{seq}}$ obtained from the encoder of the contextual module are considered as two different modalities. To enforce alignment constraints across modalities, we propose a cross-layer cross-modal alignment constraint (CAC) loss, as shown in Figure 2. The CAC loss is implemented as a knowledge distillation loss (KL-divergence here), where the entire deep contextual features and the structural features are treated as teacher and student models, respectively (the teacher model stops back-propagating gradients). A high temperature $T=8$ is adopted to “soften” the probability distribution from potential spike responses[34]. The distillation process is expressed as:
@@ 3.2.3 Cross-modal Alignment
Overall, to align the generated sequences’ probability distribution with the reference distribution, we use $\mathcal{L}_{\text{expCE}}$ to fit the probability distribution. To achieve cross-layer cross-modal alignment, we use the CAC loss to align the multi-modality features by providing long-range supervision to the structure extractor. The final objective function of MMDesign training is composed of the primary $\mathcal{L}_{\text{expCE}}$ loss and the auxiliary $\mathcal{L}_{\text{CAC}}$ loss.

### v2

# Progressive Multi-Modality Learning for Inverse Protein Folding

@@ Abstract
While deep generative models show promise for learning inverse protein folding directly from data, the lack of publicly available structure-sequence pairings limits their generalization. Previous improvements and data augmentation efforts to overcome this bottleneck have been insufficient. To further address this challenge, we propose a novel protein design paradigm called MMDesign, which leverages multi-modality transfer learning. To our knowledge, MMDesign is the first framework that combines a pretrained structural module with a pretrained contextual module, using an auto-encoder (AE) based language model to incorporate prior protein semantic knowledge. Experimental results, only training with the small dataset, demonstrate that MMDesign consistently outperforms baselines on various public benchmarks. To further assess the biological plausibility, we present systematic quantitative analysis techniques that provide interpretability and reveal more about the laws of protein design.
Index Terms:
Protein Design, Inverse Folding, Multi-Modality, Transfer Learning, Interpretability
@@
Figure 1: (a) The mainstream advanced deep protein design framework, where the GNN-based structure module represents structural features, followed by an optional contextual module, and the decoder responsible for decoding sequences. (b) The proposed novel paradigm of protein design framework consists of two pretrained modules, i.e., the pretrained structure module and the auto-encoder (AE) contextual module. Additional cross-layer cross-modal alignment aims to enhance constraints.
I Introduction
@@
De novo protein design, also known as protein inverse folding, aims to generate protein sequences based on corresponding protein coordinates. Protein design is crucial for various applications in bioengineering, including drug design and enzyme discovery, but remains challenging. Recent years have seen significant progress in protein design with the emergence of deep learning. In particular, supervised generative models have enabled more efficient and effective prediction of protein sequences[1, 2, 3, 4, 5, 6, 7, 8]. Deep generative models show the potential to learn protein design rules directly from data, making them a promising alternative to traditional approaches.
@@
The mainstream and advanced generative protein design frameworks typically follow an encoder-decoder architecture, as illustrated in Figure 1(a). A graph is used to represent the structural information in Euclidean space, and graph neural networks (GNNs) are responsible for encoding geometric information into high-dimensional space. Then the decoder recovery the protein sequences through either autoregressive networks based on self-attention[1, 4] or non-autoregressive networks[9, 10, 11]. However, the lack of massively available parallel datasets has become a major bottleneck, leading to problems such as under-training or over-fitting. Although billions of protein sequences are available, few protein structures have been experimentally determined. Hence, it is still challenging to train a practical protein design model compared to established cross-modal tasks such as vision-language works.
To address this limitation, various solutions have emerged. 1) The most common method is to focus on protein design architecture, like enhancing the structural equivalent representation, as shown in Figure 1(a). But at this stage of development, this will only marginally improve performance. 2) The most straightforward approach is to expand the dataset by generating large-scale pseudo-structures[5] sampled from advanced protein structure prediction models. However, the data construction process is tedious and time-consuming. 3) Another newly emerging framework is to introduce a pretrained structural module, which greatly reduces the reliance on supervised data while maintaining satisfactory performance [11]. Among these solutions, utilizing pretrained structural networks has become a consensus. To further enhance semantic knowledge, it is worth considering injecting the prior language knowledge of sequences since there are large-scale sequence resources. However, incorporating prior knowledge of both source (structural) and target (textual) modalities is intractable for supervised modes, particularly in protein design where there are no reference works.
Fortunately, a new paradigm in the visual-textual domain [12] enlightens us to propose a novel paradigm for generative protein design, MMDesign, which leverages multi-modality transfer learning through pretrained structural and contextual modules, as illustrated in Figure 1(b). This framework fully exploits pretraining models to overcome the data bottleneck. For the structural module, we incorporate an advanced equivariant network. For the contextual module, we introduce an auto-encoder (AE)-like manner, which can integrate comprehensive prior knowledge and assume the functions of encoding and decoding. Furthermore, we incorporate explicit cross-modal alignment constraints[13, 14, 15, 16]. This alignment can be viewed as a consistency to measure the distance between different distributions [17, 18, 12, 19, 20, 21]. Maintaining consistency across modalities allows the structural module to be constrained by long-range temporal context, as shown in Figure 1(b). The AE module here serves similar functions as consistency.
MMDesign is trained solely on the small dataset CATH [22, 1], yet extensive experiments demonstrate that it substantially outperforms the state-of-the-art baselines without relying on large-scale supervised data. However, previous studies have focused primarily on average performance, with little attention paid to explaining the generated results and their biological plausibility. We present the first systematic quantitative and comprehensive analysis techniques to assess the plausibility of the generated protein sequences and the distribution of data.
The main contributions are as follows:
$\bullet$ The proposed MMDesign innovatively introduces both prior structural and contextual semantic knowledge based on pretraining. This provides a new way of thinking to overcome the bottleneck of other cross-modal protein tasks.
$\bullet$ The introduction of explicit cross-modal alignment between the structural and contextual modality using a cross-layer consistency constraint and implicit cross-modal alignment using favourable properties of the AE mechanism.
$\bullet$ Despite only training on a small dataset, MMDesign substantially outperforms baselines and even surpasses the strong baselines trained on large datasets. More comprehensive interpretable analyses are presented for multidimensional validation and reveal the underlying protein design patterns.
@@
II Related Work
@@
Recently, generative models for inverse protein folding have emerged gradually. For instance, Structured Transformer[1] introduced the formulation of fixed-backbone design as a conditional sequence generation problem and used invariant features with GNNs. GVP-GNN [4] further improved GNNs with translation- and rotation-equivariance to enable geometric reasoning, highlighting the powerful structural geometric vector perceptron (GVP). More recently, ProteinMPNN [6] achieved higher-level sequence recovery and improved inference efficiency. Then PiFold [10] contained a novel residue featurizer and GNN layers to predict protein sequences in a one-shot way. But it struggles with longer sequences. ESM-IF[5] utilized GVP-based encoder to extract geometric features followed by a Transformer. It also employed a data augmentation strategy with nearly 12 million pseudo-protein structures, achieving an impressive native sequence recovery. However, scaling up the data by mostly three orders of magnitude significantly increases the training difficulty. Moreover, its better performance does not extend to single and short chains.
III Methods
@@
As depicted in Figure 2, MMDesign framework consists of a structural module and a contextual module sequentially. The goal of the MMDesign is to generate amino acid sequences based on the atomic coordinates (N, Cα, C) of the corresponding protein backbones. The overall training pipeline is divided into two main steps. Step 1 is to initialize the GVPConv-based structural module and pretrain the AE network for the contextual module. The parameters of the structural module are directly derived from the off-the-shelf protein structure model [5]. And step 2 is to transfer the pretrained structural module and contextual module into the MMDesign framework.
III-A Pretraining Stage
@@
Geometric Structural Module Initialization. The GVPConv structural module is a key component. GVP has been widely used in protein design[4, 5] due to its lightweight and efficient linear transformation, as well as its ability to learn rotation-equivariant transformations of vector features and rotation-invariant transformations of scalar features. In our framework, we use a variant GVP as in ESM-IF [5], which transforms the vector features into local reference frames defined for each amino acid to derive rotation-invariant features. The full graph convolution following GVP utilizes message passing to update the node embeddings in each propagation step, taking into account messages from the neighbor nodes and edges [5, 4]. To obtain a better representation of the protein structures, we initialize the GVPConv parameters as [11], where the GVPConv module is pretrained with guidance from a large-scale protein language model through contrastive alignment constraints.
@@
Contextual Transformer Pretraining. Our AE model is a generic autoregressive encoder-decoder Transformer architecture[23]. The only modification is the use of learned positional embeddings instead of classical sinusoidal positional embeddings. Shared token embedding and shared positional encoding are used since the encoder and decoder share the same amino acid tokens. To pretrain the AE solely based on the sequence data from the CATH training set, we employ a sequence-to-sequence recovery task as a pseudo-machine translation process, as shown in Figure 2. This enables the AE to learn contextual semantic knowledge. And cross-entropy loss is used to learn recovering sequences. Formally, we assume that the input protein sequence of the encoder is denoted as $S_{\text{in}}=\{s_{1},s_{2},\cdots,s_{n}\}$ with $n$ amino acids, and the generated sequence from decoder is denoted as $S_{\text{out}}=\{s^{\prime}_{1},s^{\prime}_{2},\cdots,s^{\prime}_{n}\}$, which corresponds to the probability distribution denoted as $S_{\text{logits}}$. The reference native sequence is $S_{\text{native}}=S_{\text{in}}=\{s_{1},s_{2},\cdots,s_{n}\}$. The cross-entropy loss is as follows:
@@
III-B Protein Design Framework Training
MMDesign Training Pipeline. The MMDesign framework concatenates the structural module and the contextual module sequentially, as illustrated in the lower section of Figure 2. The contextual module only preserves the encoder and decoder layers of the AE model in step 1 and discards the embedding layers. Let $X=\{x_{1},x_{2},\cdots,x_{n}\}$ denote the atomic coordinates of the protein backbone of a given structure-sequence pair, where $n$ is the length of the protein sequence, and $x_{i}\in\mathbb{R}^{3\times 3}$ represents the $i-$th 3D coordinates of the amino acid residue, consisting of N, Cα, and C atoms. The corresponding generated protein sequence is represented as $Y=\{Y_{1},Y_{2},\cdots,Y_{n}\}\in\mathbb{R}^{n}$ with the same length as $X$, where $Y_{i}$ denotes the $i-$th amino acid residue in textual form. The reference native sequence of $Y$ is $\hat{Y}=\{\hat{Y}_{1},\hat{Y}_{2},\cdots,\hat{Y}_{n}\}\in\mathbb{R}^{n}$. Firstly, the protein backbone $X$ is input to the GVPConv structural module to obtain the geometric structural features $Z_{\text{struc}}=\{z_{0},z_{1},\cdots,z_{n}\}$ with a dimension $d$. Then, the structural features $Z_{\text{struc}}$ are fed to the encoder-decoder layers of the contextual module sequentially. The features obtained by the encoder are referred to contextual features $Z_{\text{seq}}=\{z^{\prime}_{0},z^{\prime}_{1},\cdots,z^{\prime}_{n}\}$. Finally, the decoder generate the sequence distribution $D_{\text{logits}}=\{o_{0},o_{1},\cdots,o_{n}\}$ as:
$Z_{\text{struc}}=\text{GVP}(X)\in\mathbb{R}^{n\times d},$ (2)
$Z_{\text{seq}}=\textrm{Encoder}(Z_{\text{struc}})\in\mathbb{R}^{n\times d}.$ (3)
$\begin{split}D_{\text{logits}}&=\textrm{Decoder}(Z_{\text{seq}})\in\mathbb{R}^{n\times M},\\ Y&=\text{argmax}\big{(}\text{Softmax}(D_{\text{logits}})\big{)}\in\mathbb{R}^{n},\end{split}$ (4)
where $M$=20 denotes 20 types of amino acids.
Group Models Perplexity Recovery (%)
@@ III-B Protein Design Framework Training
TABLE I: Perplexity and recovery comparison. The best results are bolded, and the second best underlined. Group 1 & 3 are trained on a small-scale CATH training set only, while Group 2 is trained on a large-scale CATH+AlaphaDB training set. The Short and Single-chain are subsets of CATH test set ALL. *: evaluated on CATH 4.3; Others are evaluated on CATH 4.2.
Generation Objective. The cross-entropy loss is commonly used in protein design tasks. By minimizing the cross entropy, we can obtain an approximate distribution that better matches the target distribution. We observed that emphasizing the salient features of the output distribution can further improve the performance of our model. To achieve this, we propose an exponential cross entropy (expCE) that accentuates the spiking effect of the output distribution. The expCE is calculated as follows:
@@
Cross-modal Alignment. We view protein design as a cross-modal task, where the structural features $Z_{\text{struc}}$ obtained from the GVPConv module and the contextual features $Z_{\text{seq}}$ obtained from the encoder of the contextual module are considered as two different modalities. To enforce alignment constraints across modalities, we propose a cross-layer cross-modal alignment constraint (CAC) loss, as shown in Figure 2. The CAC loss is implemented as a knowledge distillation loss (KL-divergence here), where the entire deep contextual features and the structural features are treated as teacher and student models, respectively. A high temperature $T=8$ is adopted to ‘soften’ the probability distribution from potential spike responses. The distillation process is expressed as:
@@
Overall, $\mathcal{L}_{\text{expCE}}$ is used to to align the generation distribution with the reference distribution, and the cross-layer cross-modal alignment CAC provides long-range supervision to the structure extractor. The final training objective is composed of the primary $\mathcal{L}_{\text{expCE}}$ loss and the auxiliary $\mathcal{L}_{\text{CAC}}$ loss.